The sum of the first two terms of an exponential sequence is 135 and the sum of the third and the fourth terms is 60. Given that the common ratio is positive, calculate the common ratio and d first term
Answer by richwmiller(17219)

Tn = t * r^(n - 1)
T2 = t * r
T2+T1=135
t*r+t=135
T3 = t * r^2
T4 = t * r^3
T3+T4=60
t*r^2+t*r^3=60,
(r^3+r^2)t = 60
t*r+t=135
(r+1)*t = 135
(r^3+r^2)t = 60
(r+1)*t = 135
r = 2/3, t1 = 81